Solve the following system by the elimination method. -2x+14-8y Select the correct choice below and, if necessary, fill in the answer box to complete your choice. O A. The solution is O B. There are infinitely many solutions. (Simplify your answer. Type an ordered pair) ° C. There is no solution.
Answer #1

Answer :

Consider the given system of equation

x + 4y = 7

- 2x + 14 = 8y

Solution by elimination method :

From first equation 4y = 7 - x

Substitute 4y = 7 - x in second equation ,

- 2x + 14 = 2(7 - x)

- 2x + 14 = 14 - 2x

0 = 0

this implies that the equations are identical and hence there are infinitely many solutions